OpenStudy (anonymous):

PLZ PLZ PLZ HELP!!! MEDAL GIVEN!! How many solutions does the system of equations have?? y = 4x + 3 and 2y - 8x = 3

OpenStudy (anonymous):

So I know that you do: 2(4x + 3) - 8x = 3

OpenStudy (anonymous):

Because you plug in the y from the second equation.

OpenStudy (anonymous):

@mathman806

OpenStudy (anonymous):

a. one b. two c. infinitely many d. none

OpenStudy (anonymous):

lets get the 2y-8x=3 into y=mx+b form 2y-8x=3 add 8x 2y=8x+3 divide by 2 y=4x+6 y=4x+3 they have the same slope but they don't have the same y-intercepts so it would be none

OpenStudy (anonymous):

Ooohh. Wow thanks!! One more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

How many solutions does the system of equations have? x = -4y + 4 and 2x + 8y = 8

OpenStudy (anonymous):

well lets get them both into y=mx+b x=-4y+4 2x+8y=8 add 4y subtract 2x x+4y=4 8y=-2x+8 saubtract x divide by 8 4y=4-x y=(-1/4)x+1 divide by 4 y=1-(1/4)x y=(-1/4)x+1 same slope and same y- intercepts so infinitely many

OpenStudy (anonymous):

Wow your great help!! Thank you!!